OpenStudy (anonymous):

Suppose a bag of marbles has 4 green, 2 red, 5 yellow, 1 brown, and 7 blue marbles. What is the probability of picking a yellow marble, replacing it, and picking another yellow marble?

OpenStudy (anonymous):

i got 5/19

OpenStudy (anonymous):

Plug these into your calculator Green- 19C4 Red- 19C2 Yellow-19C5 Brown-19C1 Blue-19C7

OpenStudy (anonymous):

You have a total of 19 marbles so all of your probabilities are out of 19. As 19 is a prime number you cannot bring any of your probabilities to lower terms. Probability of a green selection is 4/19 , red is 2/19 , yellow is 5 /19 , brown is 1/19 , and finally blue which has the greatest probability of 7/19. All of these probabilities depends on the drawer of the marbles not being able to see the marbles. If you can see inside the bag then drawing the marble you want would be 100% certain lol :)

OpenStudy (anonymous):

i don't have a calculator. I can't plug it in. and so 5/19 is correct?

OpenStudy (anonymous):

yellow is 5 /19

OpenStudy (anonymous):

so yeah

OpenStudy (kropot72):

There are 19 marbles. On the first selection the probability of picking yellow is 5/19. On the second selection the probability of picking yellow is also 5/19. The outcomes of the selections are independent of each other so the combined probability is\[P(yellow/yellow)=\frac{5}{19}\times \frac{5}{19}=you\ can\ calculate\]

OpenStudy (anonymous):

exactly